10 gallons of a 15.0% alcohol solution are to be mixed with a 28.0% alcohol solution to make a 20.0% alcohol solution. How many gallons of a 28.0% alcohol must be used? How many gallons of a 20.0% alcohol solution are made?
Answer by mananth(16946)

------ Percent ---------------- quantity
Alcohol 15 ----------------10
Alcohol II 28----------------x
Total 20 ----------------10+x

15*10+28*x=20(10+x)
150+28x =200+20 x
28 x - 20 x = 200 - 150
8x=50
/8
x=6.25 gallons 28% Alcohol II
Balance 13.75 gallons 15% alcohol